SiLC beam test - CERN SPS June 2008 - HIT MODE
4 Belle modules, 40MHz, HV=80V, 6 samples
trigger: 9 mm opening, TLUControl.exe -d 10 -a 12
telescope: 0sup., new firmware
Flex_Module has only 1 working APV - control bonds suffered during transportation
Position is approximately similar to KEK Nov 2007, however it appears like JP_module is flipped (up-down) w.r.t.
to other modules
SAME AS PREVIOUS, BUT 0-SUP WITHOUT TELESCOPE
stuck after 45020 events with the repeated message
***WARNING*** FIFO3 empty, FADC 0. Event discarded, reset sent.
to be investigated.